mr. valdez is twice as old as his son. 8 years ago, he was 4 times as old as his son. what are the present ages of the father and the son?

Found 2 solutions by ankor@dixie-net.com, MathTherapy:
Answer by ankor@dixie-net.com(22740)

mr. valdez is twice as old as his son.
v = 2s
8 years ago, he was 4 times as old as his son.
v - 8 = 4(s-8)
v - 8 = 4s - 32
v = 4s - 32 + 8
v = 4s - 24
what are the present ages of the father and the son?
Replace v in the above equation, with 2s
2s = 4s - 24
2s - 4s = -24
-2s = -24
s = -24/-2
s = +12 yrs old is the son
then obviously
2(12) = 24 yrs old is Mr V
:
:
:
Check this in the statement
" 8 years ago, he was 4 times as old as his son. "
24 - 8 = 4(12 - 8)
16 = 4(4)

Answer by MathTherapy(10552)

mr. valdez is twice as old as his son. 8 years ago, he was 4 times as old as his son. what are the present ages of the father and the son?
One is 24 and the other is 12. Who makes up these problems>? I'm sure this must've happened somewhere at some time or another where a
male became a father at 12, but is that customary? I think not, so why state these things which is in most cases, unlikely? What is
WRONG with saying the 2 are brothers, sisters, cousins? Why father and son which suggests that a child at 12 years-old fathered a child?